OpenStudy (anonymous):

Solve the system using elimination by addition. x + y + z = 0 -x + y - 2z = 6 2x + y + z = 0

OpenStudy (anonymous):

If the first equation is subtracted from the third equation the result is x = 0. The equations then reduce to y + z = 0 y - 2z = 6 y + z = 0 (which is just a copy of the first equation) If we multiply the first equation by 2 and add it to the second equation we get 3y = 6 or y = 2. And that forces z = -2.